# 4DGen: Grounded 4D Content Generation with Spatial-temporal Consistency

### v1

@@ 4.3 Evaluation Metric
We use the widely adopted CLIP distance as in previous 3D content creation works [77, 63] to measure the view synthesis quality. For temporal smoothness, we employ CLIP-T as in prior works [13, 17], which is the average CLIP distance between adjacent frames. Since novel view synthesis gets more challenging as novel-view camera gets more distant, we choose to also measure CLIP-T distance at different viewpoints, not only for the frontal view but also for the back and side views, denoted as CLIP-T-f, CLIP-T-b, and CLIP-T-s, respectively.
Figure 6: Ablation studies of our proposed components. Models lacking either $\mathcal{L}_{\text{SDS}}$ or $\mathcal{L}_{\text{pseudo}}$ fail to generate high-quality details.
Figure 7: Image-to-4D Results. Our pipeline supports the Image-to-4D task by a video sequence generated by an image-to-video generative model [3].
4.4 Qualitative and Quantitative Comparisons
@@
Since the code of Consistent4D [28]is only partially released, we are unable to conduct quantitative comparisons with Consistent4D under the same settings. Our 4DGen significantly outperforms baselines in spatial and temporal consistency, evident in enhanced details and smooth transitions between frames. Furthermore, previous methods only support the generation of a fixed number of frames, which is equal to the number of input images. Our method, on the other hand, benefits from the seamless consistency priors, and supports sampling videos of arbitrary length from arbitrary viewpoints. As shown in Table. 1, Our 4DGen highly surpasses the compared methods in term of CLIP-T-f, CLIP-T-b, and CLIP-T-s and achieves competitive performances regarding CLIP metric. The better spatio-temporal continuity reflected in the CLIP-T metric aligns with observations made in the videos.
Table 1: Quantitative comparisons on the four datasets.
Dataset Metrics DreamGaussian One-2-3-45 Ours
[…]
CLIP-T-$\downarrow$ 0.0310 0.0770 0.0124
CLIP-T-$\downarrow$ 0.0463 0.0866 0.0082
CLIP-T-$\downarrow$ 0.0754 0.0807 0.0169
[…]
CLIP-T-$\downarrow$ 0.0347 0.0553 0.0095
CLIP-T-$\downarrow$ 0.0441 0.0516 0.0135
CLIP-T-$\downarrow$ 0.0788 0.0455 0.0142
[…]
CLIP-T-$\downarrow$ 0.0291 0.0936 0.0088
CLIP-T-$\downarrow$ 0.0529 0.1062 0.0118
CLIP-T-$\downarrow$ 0.0694 0.1389 0.0105
[…]
CLIP-T-$\downarrow$ 0.0264 0.0416 0.0132
CLIP-T-$\downarrow$ 0.0468 0.0593 0.0107
CLIP-T-$\downarrow$ 0.0566 0.0583 0.0235
@@
w/o $\mathcal{L}_{\text{SDS}}$ 0.3004 0.0157 0.0191 0.0724
w/o $\mathcal{L}_{\text{sync}}$ 0.3096 0.0137 0.0312 0.0729
w/o $\mathcal{L}_{\text{recon}}$ 0.2550 0.0099 0.0163 0.0143
w/o $\mathcal{L}_{\text{TV}}$ 0.2540 0.0078 0.0113 0.0144
w/o $\mathcal{L}_{\text{chamfer}}$ 0.2493 0.0096 0.0142 0.0153
Full model 0.2491 0.0095 0.0135 0.0142
4.5 Ablation Study
We conduct ablation studies to validate the effectiveness of our novel components. As shown in Tab. 2 and Fig. 6, when training without $\mathcal{L}_{\text{SDS}}$, $\mathcal{L}_{\text{recon}}$, $\mathcal{L}_{\text{sync}}$, $\mathcal{L}_{\text{TV}}$ or $\mathcal{L}_{\text{chamfer}}$, the output video suffer from distorted geometric shapes or unsatisfactory flickering artifacts. In comparison, our full model delivers the best performance.
4.6 Limitations

### v3

@@ 4.3 Evaluation Metric
We use the widely adopted CLIP distance as in previous 3D content creation works [82, 69] to measure the view synthesis quality. For temporal smoothness, we employ CLIP-T as in prior works [15, 20], which is the average CLIP distance between adjacent frames. Since novel view synthesis gets more challenging as novel-view camera gets more distant, we choose to also measure CLIP-T distance at different viewpoints, not only for the frontal view but also for the back and side views, denoted as CLIP-T-f, CLIP-T-b, and CLIP-T-s, respectively. To measure the quality of videos, we use X-CLIP [4], a constrastive pre-trained video-text model from front, back and 360 views, denoted as XCLIP-f, XCLIP-b and XCLIP-r. The textual description is human-generated.
Figure 6: Text-to-4D Comparison. Our input video sequences are generated through the process of text-to-image via Stable Diffusion [60] and image-to-video via Stable Video Diffusion [3]. Our results are more reasonable in detail and there are no noticeable artifacts. In comparison, 4D-fy results in limited motion. We provide video comparisons in the project page.
Table 1: Average results on 13 datasets.
Method DreamGaussian One-2-3-45 Consistent4D Ours
[…]
CLIP-T-$\downarrow$ 0.0275 0.0796 0.0203 0.0135
CLIP-T-$\downarrow$ 0.0647 0.0937 0.0178 0.0210
CLIP-T-$\downarrow$ 0.0491 0.0693 - 0.0129
XCLIP-$\uparrow$ 17.9058 13.8918 17.1490 18.0234
XCLIP-$\uparrow$ 16.1569 15.2904 13.8882 17.5254
XCLIP-$\uparrow$ 13.7809 15.9654 - 17.5805
4.4 Video-to-4D Qualitative and Quantitative Comparisons
@@
Our 4DGen significantly outperforms baselines in spatial and temporal consistency, evident in enhanced details and smooth transitions between frames. Furthermore, previous methods only support the generation of a fixed number of frames, which is equal to the number of input images. Our method, on the other hand, benefits from the seamless consistency priors, and supports sampling videos of arbitrary length from arbitrary viewpoints. Since we utilize the output of Consistent4D [32] as reported on its webpage, testing two metrics under identical settings is not feasible. As shown in Table. LABEL:tab:comparsion, we compare average results on 13 cases. Our 4DGen highly surpasses the compared methods across all metrics.
4.5 Image/Text-to-4D Comparisons
Our study supports input videos generated from video diffusion models, facilitating the accomplishment of image-to-4D and text-to-4D tasks. As shown in Fig. 5, we compare our image-to-4D results with Animate124 [94]. Animate124 utilizes a pretrained video diffusion model to driven object motion by video SDS loss. Despite employing textual inversion [18] to maintain identity with the reference image, it unavoidably leads to identity inconsistency. In contrast, our grounded 4D content generation faithfully generates objects in accordance with input signals.
Due to the flickering and inconsistent features in the current text-to-video model, we address text-to-4D challenges by employing a text-to-image model [60, 54] to generate the desired front image. Subsequently, we leverage an image-to-video diffusion model [3] to generate the desired video, feeding it into our pipeline to accomplish the text-to-4D task. As illustrated in Fig 6, our model achieves better details and geometry shapes. The motion in 4D-fy [6] relies on video SDS loss, inevitably introducing temporal consistency issues with flickering in video results and limiting motion variability. Our results exhibit good consistency and rich motion. We provide video comparisons in the project page.
4.6 Training Time and Memory
Our model is trained for about an hour on a scene with one RTX3090 with 15G memory. DreamGaussian [69] is a per-frame generation baseline, which requires 32 minutes(2 min per frame). Consistent4D [32] runs for 2.75 hours on a V100 with 40G memory. MAV3D [68] takes 6.5 hours on 8 A100s. 4D-fy [6] takes 23 hours on a A100. We test Animate124 [94] on a A6000, requiring approximately 4.5h. In comparison with the previous method, our model achieves fast training and requires less memory.
4.7 Ablation Study
We conduct ablation studies to validate the effectiveness of our novel components on the fish dataset. As shown in Tab. 2 and Fig. 7, when training without $\mathcal{L}_{\text{recon}}$, $\mathcal{L}_{\text{SDS}}$, $\mathcal{L}_{\text{pseudo}}$, $\mathcal{L}_{\text{TV}}$ or $\mathcal{L}_{\text{smooth}}$, the output videos suffer from distorted geometric shapes, unsatisfactory flickering artifacts or temporal choppiness. In comparison, our full model delivers the best performance.
Figure 7: Ablation studies of our proposed components. Models lacking either $\mathcal{L}_{\text{recon}}$, $\mathcal{L}_{\text{SDS}}$ or $\mathcal{L}_{\text{pseudo}}$ fail to generate high-quality details. Models without $\mathcal{L}_{\text{TV}}$ or $\mathcal{L}_{\text{smooth}}$ fail to generate outputs with smoothness. We provide video comparisons in the project page.
@@ 4.7 Ablation Study
w/o $\mathcal{L}_{\text{recon}}$ 0.2828 0.0401 0.0126 0.0401 14.8765 15.0583 13.7076
w/o $\mathcal{L}_{\text{SDS}}$ 0.3044 0.0150 0.0188 0.0994 17.5322 13.1683 13.8437
w/o $\mathcal{L}_{\text{pseudo}}$ 0.3103 0.0137 0.0286 0.0978 16.8965 14.2132 10.8373
w/o $\mathcal{L}_{\text{TV}}$ 0.2603 0.0091 0.0108 0.0158 17.2587 17.8444 16.6795
w/o $\mathcal{L}_{\text{smooth}}$ 0.2549 0.0103 0.0148 0.0205 16.3765 15.2553 16.3614
Full model 0.2491 0.0072 0.0079 0.0068 17.5926 18.4030 16.8225
4.8 Limitation